# URS: A Unified Neural Routing Solver for Cross-Problem Zero-Shot Generalization

### v1

@@ Appendix I Detailed Experimental Results
[…]
As described in Appendix E.3, the majority of VRP variants are already solved without supplying an additional problem state $C_{t}$. To test whether the remaining use of $C_{t}$ matters, we conduct an ablation in which we eliminate it entirely, enforcing $C_{t}\equiv 0$ for all problems. The results in Table 14 show that URS maintains its competitive performance even in the absence of $C_{t}$. This indicates that explicit constraint-state features are not essential for the effectiveness of URS. Interestingly, on (O)CVRPTW instances, the model without any explicit constraint features outperforms the version that supplies only capacity information. We guess that capacity and time windows information are equally important, so enforcing only capacity creates a bias toward load considerations.
@@
Table 15: Ablation of prior components in the Mixed Bias Module (MBM).
@@
[…]
To further assess the effectiveness of MBM, we conduct an ablation study that incorporates different combinations of problem-specific priors into the MBM. Owing to its flexibility, the MBM permits imposing multiple priors on the same shared attention layer. The results in Appendix K.3 underscore the complementary roles of the three prior components: (1) The transposed distance matrix $\bm{D^{\mathrm{T}}}$ captures directional asymmetry absent from the raw distance matrix, and its inclusion markedly stabilizes optimization on asymmetric variants; (2) Removing the pickup–delivery relation matrix $\bm{R}$ degrades performance on PDTSPs, confirming that explicit relational coupling is indispensable for these instances; (3) When all priors are fused, MBM can simultaneously learn the geometric and relational biases inherent in various problems, yielding the best overall performance.
@@
Table 16: Effect of different decoder architectures on URS cross‑problem performance
@@
We adopt an AM (Kool et al., 2019) as our basic encoder-decoder model. Multiple decoder refinements have been proposed around AM (Kwon et al., 2020; Huang et al., 2025; Zhou et al., 2024a). To assess how decoder structure impacts multi-task performance, we additionally train two URS variants whose decoders follow POMO and ReLD, denoted URS-POMO and URS-ReLD. For stronger and fairer baselines, we apply Distance-aware Attention Reshaping (DAR) (Wang et al., 2025b) to both, and all decoder parameters are generated by a hypernetwork conditioned on the problem representation $\bm{\lambda}$ (see Appendix E.2). As shown in Table 16, URS-POMO and URS-ReLD still exhibit strong cross-problem robustness, while replacing the decoder with ICAM (Zhou et al., 2024a) yields a further overall improvement.
@@
HGS-PyVRP (Wouda et al., 2024) Code https://github.com/PyVRP/PyVRP MIT License
@@
OR-Tools (Perron & Furnon, 2023) Code https://github.com/google/or-tools Apache-2.0 License
@@
ICAM (Zhou et al., 2024a) Code https://github.com/CIAM-Group/ICAM MIT License
@@
MTPOMO (Liu et al., 2024a) Code https://github.com/FeiLiu36/MTNCO MIT License
MVMoE (Zhou et al., 2024b) Code https://github.com/RoyalSkye/Routing-MVMoE MIT License
RouteFinder (Berto et al., 2024) Code https://github.com/ai4co/routefinder MIT License
@@
We list the used existing codes and datasets in Table 17, and all of them are open-sourced resources for academic usage.
Appendix M LLM usage
During the preparation of this manuscript, we utilized a Large Language Model (LLM) as a writing assistance tool. The primary purpose of its use was to improve the linguistic quality of the paper. Specifically, the LLM is used to polish the writing for enhanced clarity and fluency and to identify and correct grammatical, spelling, and punctuation errors.
We confirm that all core research ideas, study design, data analysis, and conclusions are independently developed by the authors. The LLM does not contribute to the research ideation, methodology, or substantive academic content of the paper. Its function was strictly confined to language editing and proofreading to ensure the manuscript’s professionalism and readability.

### v2

@@ Appendix I Generalization to Unseen VRPs
w/o $C_{t}$ 2.09% 0.45% 0.60% 1.20% 4.59% 4.03% 2.70% 5.71% 2.64% 4.02% 4.62% 2.97%
w/ $C_{t}$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06% 1.81% 6.13% 1.46% 3.24% 5.07% 2.74%
As described in Appendix E.3, the majority of VRP variants are already solved without supplying an additional problem state $C_{t}$. To test whether the remaining use of $C_{t}$ matters, we conduct an ablation in which we eliminate it entirely, enforcing $C_{t}\equiv 0$ for all problems. The results in Table 22 show that URS maintains its competitive performance even in the absence of $C_{t}$. This indicates that explicit constraint-state features are not essential to URS effectiveness. Interestingly, on (O)CVRPTW instances, the model without any explicit constraint features outperforms the version that supplies only capacity information. We guess that capacity and time windows information are equally important, so enforcing only capacity creates a bias toward load considerations.
@@
Table 23: Ablation of prior components in the MBM.
@@
$\checkmark$ $\times$ $\times$ 16.64% 0.44% 0.25% 0.67% 7.33% 16.62% 2.42% 6.44% 3.96% 9.44% 5.47% 6.33%
$\checkmark$ $\checkmark$ $\times$ 2.09% 0.45% 0.48% 0.82% 7.66% 9.11% 2.52% 6.89% 2.68% 5.70% 6.36% 4.07%
$\checkmark$ $\times$ $\checkmark$ 16.60% 0.46% 0.49% 0.94% 4.49% 26.22% 4.16% 6.57% 3.07% 12.23% 5.53% 7.34%
$\checkmark$ $\checkmark$ $\checkmark$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06% 1.81% 6.13% 1.46% 3.24% 5.07% 2.74%
To further assess the effectiveness of MBM, we conduct an ablation study that examines different combinations of problem-specific priors within MBM. Owing to its flexibility, the MBM allows multiple priors to be imposed on the same shared attention layer. The results in Table 23 underscore the complementary roles of the three prior components: (1) The transposed distance matrix $\bm{D^{\mathrm{T}}}$ captures directional asymmetry absent from the raw distance matrix, and its inclusion markedly stabilizes optimization on asymmetric variants; (2) Removing the pickup-delivery relation matrix $\bm{R}$ degrades performance on PDTSPs, confirming that explicit relational coupling is indispensable for these instances; (3) When all priors are fused, MBM can simultaneously learn the geometric and relational biases inherent in various problems, yielding the best overall performance.
@@
Table 24: Effect of different decoder architectures on URS cross-problem performance.
@@
We adopt an AM (Kool et al., 2019) as our basic encoder-decoder model. Multiple decoder refinements have been proposed around AM (Kwon et al., 2020; Huang et al., 2025; Zhou et al., 2026). To assess how decoder structure impacts multi-task performance, we additionally train two URS variants whose decoders follow POMO and ReLD, denoted URS-POMO and URS-ReLD. For stronger and fairer baselines, we apply Distance-aware Attention Reshaping (DAR) (Wang et al., 2025b) to both, and all decoder parameters are generated by a hypernetwork conditioned on the problem representation $\bm{\lambda}$ (see Appendix E.2). As shown in Table 24, URS-POMO and URS-ReLD still exhibit strong cross-problem robustness, while replacing the decoder with ICAM (Zhou et al., 2026) yields a further overall improvement.
Furthermore, the results in Table 24 reveal a notable optimization challenge. By incorporating an additional Feed-Forward Network layer compared to POMO, the ReLD decoder forces the hypernetwork to map to a larger parameter space, thereby increasing its learning difficulty and degrading performance. This optimization difficulty prevents URS-ReLD from matching the performance of the original ReLD-MTL, where weights are learned directly rather than being dynamically generated. Given the outstanding performance of ReLD and other heavy-decoder paradigms in single-task optimization (Luo et al., 2023; Drakulic et al., 2023), exploring how to adapt problem-conditioned generation mechanisms for heavier decoders is a highly promising direction.
K.5 Effects of MBM & Parameter Generator on Zero-shot Generalization
Table 25: Ablation of MBM, $\mathrm{WEIGHT}(\bm{\lambda})$, and $\mathrm{BIAS}(\bm{\lambda})$ on cross-problem zero-shot generalization. The symbol (*) indicates the seen problems in training.
MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ ATSP (*) TSP (*) OP (*) PCTSP (*) PDTSP (*) ACVRP (*)
$\times$ $\checkmark$ $\checkmark$ 16.64% 0.44% 0.25% 0.67% 7.33% 16.62%
$\checkmark$ $\times$ $\times$ 4.24% 1.10% 1.04% 1.42% 6.38% 4.07%
$\checkmark$ $\times$ $\checkmark$ 3.83% 1.16% 1.09% 1.52% 6.65% 4.27%
$\checkmark$ $\checkmark$ $\times$ 2.79% 1.58% 1.39% 1.92% 7.05% 3.39%
$\checkmark$ $\checkmark$ $\checkmark$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06%
MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ CVRP (*) CVRPTW (*) CVRPB (*) OCVRP (*) OCVRPTW (*) CVRPL
$\times$ $\checkmark$ $\checkmark$ 2.42% 6.44% 3.96% 9.44% 5.47% 0.98%
$\checkmark$ $\times$ $\times$ 6.23% 7.22% 5.00% 3.99% 5.41% 2.92%
$\checkmark$ $\times$ $\checkmark$ 4.94% 6.08% 3.00% 4.61% 5.21% 4.79%
$\checkmark$ $\checkmark$ $\times$ 2.90% 6.60% 3.12% 4.47% 4.76% 1.46%
$\checkmark$ $\checkmark$ $\checkmark$ 1.81% 6.13% 1.46% 3.24% 5.07% 0.43%
MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
$\times$ $\checkmark$ $\checkmark$ 15.95% 4.18% 2.89% 18.78% 10.93% 9.51%
$\checkmark$ $\times$ $\times$ 9.54% 5.30% 2.73% 14.37% 9.80% 4.64%
$\checkmark$ $\times$ $\checkmark$ 9.58% 3.25% 3.07% 14.16% 9.91% 3.88%
$\checkmark$ $\checkmark$ $\times$ 9.92% 3.34% 3.08% 14.02% 9.92% 4.52%
$\checkmark$ $\checkmark$ $\checkmark$ 9.35% 1.65% 2.67% 13.77% 9.18% 3.22%
MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW Avg.gap Best Sol.
$\times$ $\checkmark$ $\checkmark$ 10.74% 16.14% 5.57% 18.80% 8.37% 3/22
$\checkmark$ $\times$ $\times$ 9.67% 10.67% 5.23% 14.49% 6.16% 0/22
$\checkmark$ $\times$ $\checkmark$ 9.63% 10.42% 5.51% 13.99% 5.93% 1/22
$\checkmark$ $\checkmark$ $\times$ 9.79% 10.02% 5.81% 13.98% 5.72% 1/22
$\checkmark$ $\checkmark$ $\checkmark$ 8.94% 9.47% 5.12% 13.72% 4.89% 17/22
To empirically evaluate the necessity of these components, we have conducted tests on the 11 seen problems as well as 11 additional unseen CVRP variants widely investigated in current multi-task research (e.g., MVMoE (Zhou et al., 2024)). The results are presented in Table 25. We can observe that removing MBM leads to a significant performance drop on problems with complex geometric properties, such as ATSP, ACVRP, and PDTSP, which confirms MBM is critical for perceiving specific geometric biases. In addition, removing $\mathrm{WEIGHT}(\bm{\lambda})$ or $\mathrm{BIAS}(\bm{\lambda})$ also results in performance degradation across most tasks. In terms of overall performance, URS significantly outperforms any model where a single module is removed. The full URS model achieves the lowest average gap and provides the best solution on the majority of tasks (17/22). This demonstrates that the synergy of these components is essential for robust cross-problem generalization.
K.6 Effects of Each UDR Component
To clarify the independent contribution of each UDR component, we conduct ablations by removing each component in turn. As shown in Table 26, the absence of any single component leads to notable performance degradation. A complete UDR achieves the lowest average gap and provides the best solution on the majority of tasks (15/22), affirming the necessity of our UDR design for robust zero-shot generalization.
Table 26: The ablation study of each UDR component. The symbol (*) indicates the seen problems in training. $\bm{\rho}$, $\bm{\omega}$, and $\bm{\xi}$ represent positional identifier, unified attribute set, and node-type indicator, respectively. Since removing positional identifiers prevents the use of instance augmentation, we have excluded it from all results to ensure consistency in the inference strategy.
$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ ATSP (*) TSP (*) OP (*) PCTSP (*) PDTSP (*) ACVRP (*)
$\times$ $\checkmark$ $\checkmark$ 7.64% 1.76% 1.17% 1.96% 8.54% 7.16%
$\checkmark$ $\times$ $\checkmark$ 10.07% 1.98% 2.70% 7.05% 13.50% 13.36%
$\checkmark$ $\checkmark$ $\times$ 5.43% 1.02% 1.08% 1.86% 12.72% 5.81%
$\checkmark$ $\checkmark$ $\checkmark$ 5.51% 1.17% 0.96% 1.79% 7.71% 6.11%
$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ CVRP (*) CVRPTW (*) CVRPB (*) OCVRP (*) OCVRPTW (*) CVRPL
$\times$ $\checkmark$ $\checkmark$ 2.68% 6.88% 2.84% 4.64% 5.91% 1.29%
$\checkmark$ $\times$ $\checkmark$ 4.15% 23.71% 5.24% 7.10% 19.10% 2.70%
$\checkmark$ $\checkmark$ $\times$ 4.72% 7.67% 2.83% 9.64% 6.96% 3.33%
$\checkmark$ $\checkmark$ $\checkmark$ 2.62% 6.80% 2.69% 4.54% 6.00% 1.23%
$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
$\times$ $\checkmark$ $\checkmark$ 11.87% 3.12% 3.34% 34.34% 44.03% 4.63%
$\checkmark$ $\times$ $\checkmark$ 15.02% 5.47% 19.53% 30.88% 29.64% 6.98%
$\checkmark$ $\checkmark$ $\times$ 26.06% 2.69% 4.13% 34.62% 26.55% 9.68%
$\checkmark$ $\checkmark$ $\checkmark$ 11.54% 3.03% 3.38% 15.47% 10.70% 4.55%
$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW Avg.gap Best Sol.
$\times$ $\checkmark$ $\checkmark$ 43.99% 11.99% 6.02% 34.66% 11.38% 3/22
$\checkmark$ $\times$ $\checkmark$ 29.34% 14.84% 19.22% 31.02% 14.21% 0/22
$\checkmark$ $\checkmark$ $\times$ 28.60% 25.96% 7.00% 35.93% 12.01% 4/22
$\checkmark$ $\checkmark$ $\checkmark$ 10.51% 11.68% 6.08% 15.38% 6.34% 15/22
K.7 Hypernetwork vs. Adapter
We train URS-Adapter, a URS variant that adopts a distinct decoder for each of the 11 variants, and compare it with the original URS using a hypernetwork on 11 seen problems (see Table 27) and large-scale CVRP variants (see Table 28). URS consistently outperforms URS-Adapter across 11 seen problems, achieving higher parameter efficiency. This performance advantage holds consistently on large-scale CVRP variants, which further solidifies our conclusion.
Table 27: Performance comparison between URS-Hypernet and URS-Adapter on 11 seen VRP variants.
Method OP PCTSP PDTSP ACVRP
Gap Time Gap Time Gap Time Gap Time
URS-Adapter 0.88% 4s 1.47% 4s 5.69% 4s 3.83% 1.5m
URS 0.45% 4s 1.06% 4s 4.98% 4s 3.06% 1.5m
Method CVRPTW CVRPB OCVRP OCVRPTW
Gap Time Gap Time Gap Time Gap Time
URS-Adapter 6.34% 8s 3.21% 5s 4.06% 6s 4.94% 7s
URS 6.13% 8s 1.46% 5s 3.24% 6s 5.07% 7s
Method ATSP TSP CVRP Average Best
Gap Time Gap Time Gap Time Gap Solution
URS-Adapter 2.97% 1.2m 0.82% 4s 2.85% 6s 3.37% 1/11
URS 2.26% 1.2m 0.57% 4s 1.81% 6s 2.74% 10/11
Table 28: Generalization comparison between URS and URS-Adapter across diverse large-scale CVRP variants.
Problem Method N = 1000 N = 2000 N = 3000 N = 4000 N = 5000
Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time
CVRPTW PyVRP 159.35(0.00%) 20m 337.15(0.00%) 40m 516.97(0.00%) 1h 618.58(0.00%) 2.7h 787.25(0.00%) 3.3h
URS-Adapter 173.68(8.99%) 1.3m 367.22(8.92%) 10.8m 564.55(9.21%) 30.8m 677.91(9.59%) 1.2h 860.63(9.32%) 2.3h
URS 172.58(8.30%) 1.3m 365.27(8.34%) 11.1m 559.03(8.14%) 30.1m 673.63(8.90%) 1.2h 853.79(8.45%) 2.3h
CVRPB PyVRP 36.52(0.00%) 20m 51.80(0.00%) 40m 75.12(0.00%) 1h 93.52(0.00%) 2.7h 112.21(0.00%) 3.3h
URS-Adapter 42.71(16.96%) 43s 62.73(21.11%) 5.5m 85.72(14.11%) 17.5m 107.17(14.59%) 41.9m 127.93(14.01%) 1.3h
URS 34.86(-4.54%) 40s 50.39(-2.72%) 5.1m 72.01(-4.15%) 17.3m 91.74(-1.91%) 41.2m 111.93(-0.24%) 1.3h
OCVRPTW PyVRP 90.91(0.00%) 20m 164.00(0.00%) 40m 224.16(0.00%) 1h 299.45(0.00%) 2.7h 367.86(0.00%) 3.3h
URS-Adapter 105.96(16.56%) 1.1m 203.52(24.10%) 9.9m 287.70(28.35%) 28.6m 391.06(30.59%) 1h 483.19(31.35%) 1.9h
URS 98.76(8.63%) 1m 178.07(8.58%) 7.6m 243.88(8.80%) 25.4m 325.64(8.74%) 1h 398.72(8.39%) 1.9h
@@
PyVRP (Wouda et al., 2024) Code https://github.com/PyVRP/PyVRP MIT License
@@
OR-Tools (Perron and Furnon, 2023) Code https://github.com/google/or-tools Apache-2.0 License
@@
ICAM (Zhou et al., 2026) Code https://github.com/CIAM-Group/ICAM MIT License
ELG (Gao et al., 2024) Code https://github.com/gaocrr/ELG MIT License
@@
MTPOMO (Liu et al., 2024) Code https://github.com/FeiLiu36/MTNCO MIT License
MVMoE (Zhou et al., 2024) Code https://github.com/RoyalSkye/Routing-MVMoE MIT License
RouteFinder (Berto et al., 2025) Code https://github.com/ai4co/routefinder MIT License
@@
CVRPLib Set-AGS (Arnold et al., 2019) Dataset http://vrp.galgos.inf.puc-rio.br/index.php/en/ Available for academic research use
We list the existing codes and datasets in Table 29, all of which are open-source resources for academic use.